You are given the following information, find the Z-statistic: (95% level of confidence) based on the Z-test
Xbar = 980
U = 1000
Sample stdev = 50
N = 36


What is the underlying difference between the T-test and the Z-test?


Answer by stanbon(75887)

You are given the following information, find the Z-statistic: (95% level of confidence) based on the Z-test
Xbar = 980
U = 1000
Sample stdev = 50
N = 36
------------
z(980) = (980-1000)/50 = -2/5
--------------------------------------
What is the underlying difference between the T-test and the Z-test?
In general, the T-test values change depending on sample size.
---------------------
Cheers,
Stan H.